OpenStudy (anonymous):

I really need help. A shopper used some 30-cents off coupons and some 20-cents off coupons at the supermarket to get a reduction of $3.30 from her grocery bill. If she used a total of 14 coupons, how many of the 30-cents-off coupons did she use?

OpenStudy (saifoo.khan):

yay!

OpenStudy (saifoo.khan):

i got yr ans

OpenStudy (saifoo.khan):

0.3(5) + 0.2(9) 3.3

OpenStudy (saifoo.khan):

so he uses 5 coupons of 30 cents

OpenStudy (anonymous):

wow! Thank you can you show me all your work this? so I can try to see how your broke it down

OpenStudy (saifoo.khan):

lol, well i did it mentally.

OpenStudy (anonymous):

its been 25 years since I did this stuff. I am so lost

OpenStudy (saifoo.khan):

lol, it happens sometimes.

OpenStudy (anonymous):

Thanks so much, been working on this hopemwork assignment for hours

OpenStudy (saifoo.khan):

i workd for minutes! ^.^

OpenStudy (saifoo.khan):

3-4mins

OpenStudy (anonymous):

still can you break it down for me sowwy

OpenStudy (saifoo.khan):

umm, i cant. idk how i did it.. i just usd my calc & did it mentally.

OpenStudy (anonymous):

I just the actual first step

OpenStudy (saifoo.khan):

radar will help u

OpenStudy (radar):

I will make an attempt to show you the steps. first I have to read the problem and see what is there. Fact: There were 14 coupons used. Fact 2: Only 20 cent and 30 cent coupons were use Fact 3 the coupons added up to a savings of $3.30 savings. I think those are the only facts we need to concern ourselves with/

OpenStudy (radar):

Let x = the number of 20 cent coupons Let y = the number of 30 cent coupons Fact: x+y =14. Do you follow so far?

OpenStudy (anonymous):

yes

OpenStudy (radar):

Good lets get to the savings! 20x + 30y =330 cents that tells us about the savings,,,,,,,it could also be written as: .20x +.30y=$3.30 Is that clear?

OpenStudy (anonymous):

yes]

OpenStudy (radar):

Now lets take these two equations with only two variaBLEs (x and y) and find out how many of each coupon she used: x+y=14 and .2x +.3y=3.30 Which one do you want to solve for first?

OpenStudy (radar):

x or y ??

OpenStudy (anonymous):

y

OpenStudy (radar):

I want to solve for y firs the number of .30 coupons. Using the elimination method I take the first equation and multiply all terms by .2 .2(x+y=14) .2x +.2y=2.8 I will then subtract the second equation form this -(.2x+.3y=3.3) ----------------- 0x -.1y =-.5 Note how the x was eliminated whoooosh it is gone. we solve for y. -.1y=-.5 multiply thru by 10 -y=-5 y=5 Just as the previous poster said. There were 5 30 cent coupons. Did you follow the solution part?

OpenStudy (radar):

did you understand how -y=-5 became y=5 ??

OpenStudy (anonymous):

hold on

OpenStudy (anonymous):

hold on

OpenStudy (radar):

to find x just plug the 5 for y in x+y=14 or x+5=14 solve for x

OpenStudy (radar):

I'm a holdin

OpenStudy (anonymous):

I GOT IT : )

OpenStudy (radar):

Great. Good luck.

OpenStudy (anonymous):

Thank you so much

OpenStudy (radar):

You are welcome, glad I helped.